OpenStudy (anonymous):

Let i0=1.0*10^-12 watt per square meter be the energy intensity of the softest sound detectable by the human ear. Let L(i)=10*log(i/i0) be the formula for the loudness of a sound, in decibels(dB), whose energy intensity is i. Find the decibel measure of the sound whose energy intensity is 8*10^-4 watt per square meter. A.80.0 dB B.89.0 dB C.98.0 dB D.107.0 dB E.116.0 dB

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Can I ask how you found that?

OpenStudy (anonymous):

L= 10log (i/i0) i = 8 * 10^-4 i/i0 = 8*10^8

OpenStudy (anonymous):

right

OpenStudy (anonymous):

log (8*10^8) = log 8 + log 10^8 = 0.903 + 8

OpenStudy (anonymous):

that into ten is 89

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

ne more?

OpenStudy (anonymous):

Find the logarithm without using a calculator: log (radical) 100000000000 =

OpenStudy (anonymous):

wts radical?

OpenStudy (anonymous):

if its log to the base 10 then count the no of zeros and thats d answer